2/3 correct answers are required for credit

Consider the insoluble compound zinc sulfide , ZnS. The zinc ion also forms a complex with hydroxide ions. Write a balanced net ionic equation to show why the solubility of ZnS (s) increases in the presence of hydroxide ions and calculate the equilibrium constant for this reaction. For Zn(OH)42. , K,-29x1015 . Use the pull-down boxes to specify states such as (aq) or (s). 1 1
Answer #1

Answers are shown in red.

(I did not find the exact value of Ksp of zinc cyanide, so if you have a reliable source, please substitute your value in part 3).

1 ZnS (s) H2O Ksp = 1.6x1024 z faq) + 5² (ar) Zn3+ (ar) + 4 ÕH (aq) = Zn(OH)2 (aq) Ke = 2.9 x 105 Zus(s) + 4 ÕH (aq) = Zn (03) H₂O Zu (CN)₂ (s) Zez+ aq) + 2 TN (aq) Ksp = 8.0x 1612 zu2+ (aq) + 4 N Hz (aq? = Zn (NH3)42* (aq) Kq 2.9 x 10 09 Zn(ON),(s)